How far can a car going 45mph travel in 40 min. Please show your. My book says 30 miles but I can't figure out how they did it. Is it d=r/t Please help, thank you Sophie age 12.
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

Divide 40 min by 60 min =.6666666 hr
45*.666666=30 mi

Answer by stanbon(75887)

How far can a car going 45mph travel in 40 min. Please show your. My book says 30 miles but I can't figure out how they did it. Is it d=r/t Please help, thank you Sophie age 12.
-----
The formula that relates distance, time, and rate is d = rt, not d = r/t.
----------------------------
Your problem:
rate = 45 mph ; time = (40/60) = (2/3) hr.
--------
Equation:
Since d = rt you get:
distance = 45*(2/3) miles
distance = 15*2 miles
distance = 30 miles
============================
Cheers,
Stan H.